ADNAN MUHAMMED

Grade 12,

A railway track (made of iron) is laid in winter when the average temperature is 18°C. The track consists of sections of 12"0 m placed one after the other. How much gap should be left between two such sections so that there is no compression during summer when the maximum temperature goes to 48°C ? Coefficient of linear expansion of iron = 11 x 10 6/°C.

A railway track (made of iron) is laid in winter when the average temperature is 18°C. The track consists of sections of 12"0 m placed one after the other. How much gap should be left between two such sections so that there is no compression during summer when the maximum temperature goes to 48°C ? Coefficient of linear expansion of iron = 11 x 10 6/°C.

 
 
 

Grade:11

1 Answers

Gaurav
6 years ago
  • As iron will expand in summer so as to avoid compression between tracks
  • Gap between two tracks=2l-2l0=2l0+2l0(\Delta \alpha​)\Deltat -2l0=2*11*10-6*(48-18)=7.92*10-4m